# Bias–variance Tradeoff in Tensor Estimation

### v1

@@ Appendix A Proof of Theorem 1
$=\left\|\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{*}\otimes U_{3}^{*})_{\perp}\right\|_{\mathrm{F}}.$
@@
$\|B-B_{r_{1}}\|_{F}=$ $\|\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}-\{\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}\}_{r_{1}}\|_{\mathrm{F}}$
@@
Appendix B Deviation Bounds
Throughout this appendix we work with centered sub-Gaussian random variables with parameter $\kappa^{2}$, and we assume (without loss of generality) that each such variable $X$ satisfies $\mathbb{E}X^{2}=\kappa^{2}$. Any other case can be handled with additional constants in the bounds.
Lemma 6 (Theorem 4.4.3 in Vershynin, (2018)).
Then there exist some universal constant $C>0$, such that
$\|Z\|\leq C\kappa\left(\sqrt{m}+\sqrt{n}\right)$
with probability at least $1-\exp(-(m+n)))$.
Lemma 7.
Then there exist some universal constants $C,c>0$, such that
with probability at least $1-\exp(-c\sum_{k=1}^{3}p_{k}r_{k})$.
Proof.
This directly follows from Lemma E.5 in Han et al., (2022). ∎
Lemma 8.
Proof.
Lemma 9.
Then there exists absolute positive constants $C_{1},C_{2}$ and $c$ such that
$\leq C_{2}\exp\left(-c(p_{1}+p_{2}+p_{3})\right),$
Proof.
It follows from Lemma 10 that
$\{A\in\mathbb{R}^{p_{2}\times r_{2}}:\|A\|\leq 1\}$
$\{B\in\mathbb{R}^{p_{3}\times r_{3}}:\|B\|\leq 1\}$
@@ Appendix B Deviation Bounds
Note that $\|V_{2}-\widetilde{V}_{2}\|\leq 1/4$. So
Similarly
In addition,
So for any $V_{2}$ and $V_{3}$,
or simply
Step 3. By Proposition 8 in Pajor, (1998), the cardinality of $\mathcal{N}_{p_{2},r_{2}}(\epsilon)$ is bounded $(\frac{C}{\epsilon})^{p_{2}r_{2}}$, and $\mathcal{N}_{p_{3},r_{3}}(\epsilon)$ is bounded $(\frac{C}{\epsilon})^{p_{3}r_{3}}$. Therefore
Lemma 10.
Proof.
$\|a-\widetilde{a}\|_{2}\leq\epsilon.$
Denote the random quantity
Then
[…]
Note that $\|a-\widetilde{a}\|\leq\frac{1}{4}$. So
Similarly
$|\widetilde{a}^{\top}AZB(\widetilde{b}-b)|\leq\frac{\psi}{4}.$
In addition,
So for any $a$ and $b$,
or simply
Step 4. By Vershynin, (2018), the cardinality of $\mathcal{N}_{p}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{p}$, and the cardinality of $\mathcal{N}_{q}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{q}$. Therefore
Lemma 11.
where $C$ and $C_{1}$ are positive constants.
Proof.
For any non-random $u,v\in\mathbb{R}^{m}$ , it follows that
$\|a-\widetilde{a}\|_{2}\leq\epsilon.$
Denote the random quantity
Then
[…]
Note that $\|a-\widetilde{a}\|_{2}\leq\frac{1}{4}$. So
Similarly
In addition,
So for any $a$ and $b$,
or simply
Step 2. By Vershynin, (2018), the cardinality of $\mathcal{N}_{p}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{p}$, and the cardinality of $\mathcal{N}_{q}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{q}$, for a positive constant $C$. Therefore
Lemma 12.
$\|u^{\top}Zv\|_{\psi_{2}}\leq\kappa\|u\|_{2}\|v\|_{2}.$
Then with probability at least $1-\exp(-c(m+n))$, it holds that
where $c,C$ and $C_{1}$ are positive constants.
Proof.
By assumption,
The rest of the proof is similar and simpler than Lemma 10 and is omitted.
B.1 SVD for Unbalanced Matrices
@@
where the last inequality uses $\frac{\sigma_{R}^{2}(X)+n{\kappa}^{2}}{n\,\kappa^{2}}\geq 1$. Thus, combining (B.1), (35), (36) and (37) with $x=\kappa\sqrt{\frac{2C}{c}m\,}$, we have
@@
It follows from (B.1), (33) and (38) that by the union bound,
@@
Suppose the conditions of Theorem 1 hold, in particular the condition in Equation (2) for each mode-$k$ unfolding $\mathcal{M}_{k}(X^{*})$ with target rank $r_{k}$. Let $U_{k}^{(0)}$ be the matrix of top $r_{k}$ left singular vectors of the mode-$k$ unfolding $Y^{(k)}:=\mathcal{M}_{k}(Y)$. Then for each $k\in\{1,2,3\}$, with probability at least $1-C_{1}\exp(-C_{2}p_{k})$,
@@
Apply Lemma 13 to $Y^{(k)}=\mathcal{M}_{k}(Y)\in\mathbb{R}^{p_{k}\times p_{\neg k}}$ with signal $X^{(k)}=\mathcal{M}_{k}(X^{*})$ and noise $Z^{(k)}=\mathcal{M}_{k}(Z)$, where $p_{\neg k}:=\prod_{j\neq k}p_{j}$. Use rank $r=r_{k}$ and identify $n:=p_{k}$ (rows) and $m:=p_{\neg k}$ (columns). The lemma (applied to left singular vectors, or equivalently to the transpose) gives
$\bigl\|\sin\Theta(U_{k}^{(0)},U_{k}^{*})\bigr\|^{2}\;\leq\;C\left\{\frac{p_{k}\,\kappa^{2}}{\Delta_{k}^{2}}+\frac{\kappa^{4}\,p_{k}\,p_{\neg k}}{\Delta_{k}^{4}}\right\},$
where $\Delta_{k}:=\sigma_{r_{k}}\!\bigl(\mathcal{M}_{k}(X^{*})\bigr)-\sigma_{r_{k}+1}\!\bigl(\mathcal{M}_{k}(X^{*})\bigr)$ is the mode-$k$ spectral gap.
@@
$\Delta_{k}^{2}\;\geq\;C_{\mathrm{gap}}\kappa^{2}\!\left(\sqrt{p_{k}\,p_{\neg k}\,r_{\max}}+r_{\max}\sum_{j=1}^{3}p_{j}\right).$
@@
Appendix C Matrix Perturbation Bounds
@@
where $B_{(R)}$ denote the rank-$R$ truncated SVD of $B$, this is, if $B=U\Sigma V^{\top}$ then $B_{(R)}:=U_{(R)}\Sigma_{(R)}V_{(R)}^{\top}$.
@@
where the first inequality follows from the triangle inequality, and second inequality follows from Weyl’s inequality (Weyl,, 1912), i.e. $|\sigma_{j}(A)-\sigma_{j}(B)|\leq\|A-B\|$ for all $1\leq j\leq n$, as well as the fact that
@@
where the last inequality follows from Theorem 23. It then follows from (C),
@@
where the first two inequalities follow from the same arguments as in (C). Consequently,
@@
Lemma 21 (Proposition 1 of Cai and Zhang, (2018)).
@@
Theorem 23 (Mirsky’s singular value inequality in Mirsky, (1960)).
@@
Lemma 25 (Ky Fan-type Inequality for Sums of Matrices).
Proof.
For a symmetric matrix $M\in\mathbb{R}^{n\times n}$, by Ky Fan’s maximum principle (see e.g. II.1.13 in Bhatia, (2013)), for any $1\leq k\leq n$,
Therefore
and so
Then

### v3

@@ Appendix B Proofs in Section 3
$\leq\left\|\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{*}\otimes U_{3}^{*})_{\perp}\right\|_{\mathrm{F}}.$
@@
$\|B-B_{r_{1}}\|_{\mathrm{F}}=$ $\|\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}-\{\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}\}_{r_{1}}\|_{\mathrm{F}}$
@@
Proof of Corollary 3.
It suffices to observe that for $k=1,2,3$
$\|X^{*}\|_{\mathrm{F}}^{2}\geq r_{k}\sigma_{r_{k}}^{2}(\mathcal{M}_{k}(X^{*})).$
The desired follows from the assumption that $r_{k}\leq\sqrt{p_{\min}}$ and (7). ∎
Proof of Theorem 4.
Step 1. We follow a similar strategy as in the proof of Theorem 2 to show
$\inf_{\widehat{X}}\sup_{X^{\ast}\in\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{\ast}\bigr\|_{\mathrm{F}}^{2}\$$\geq\ c(\kappa^{2}r_{1}r_{2}r_{3}+\xi^{2}).$ (27)
Let
$r_{1}^{\prime}=\lfloor\xi^{2}/(r_{2}r_{3}\kappa^{2})\rfloor.$
Consider the set of tensors
$\mathcal{B}=\{B\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}:\|B\|_{\infty}\leq\kappa^{2},B_{i,j,k}=0\text{ if }i>r_{1}+r_{1}^{\prime}\text{ or }j>r_{2}\text{ or }k>r_{3}\}.$
In this step, we show that
$\mathcal{B}\subset\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}.$ (28)
To see this, let $B\in\mathcal{B}$, and denote $B^{\prime}$ to be the matrix such that
$B^{\prime}_{i,j,k}=\begin{cases}B_{i,j,k}&\text{if }i\leq r_{1}\text{ and }j\leq r_{2}\text{ and }k\leq r_{3};\\ 0&\text{otherwise}.\end{cases}$
Then $B^{\prime}\in\mathcal{T}_{(r_{1},r_{2},r_{3})}$ and
$\|B-B^{\prime}\|_{\mathrm{F}}^{2}\leq r^{\prime}_{1}r_{2}r_{3}\|B\|_{\infty}=r^{\prime}_{1}r_{2}r_{3}\kappa^{2}\leq\xi^{2}.$
Therefore $\mathcal{B}\subset\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}$. So
$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\ \geq\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{B}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}.$ (29)
Step 1.1. Note that the tensors in $\mathcal{B}$ are having at most $(r_{1}+r_{1}^{\prime})r_{2}r_{3}$ nonzero entries, the set of matrices $\mathcal{B}$ can be viewed as vectors in $\mathbb{R}^{d}$, with $d=(r_{1}+r_{1}^{\prime})r_{2}r_{3}$. Denote
$\mathcal{V}=\{V\in\mathbb{R}^{d},\|V\|_{\infty}\leq\kappa\}.$
@@
$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{B}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}=\inf_{\widehat{V}}\sup_{V\in\mathcal{V}}\mathbb{E}\|\widehat{V}-V\|^{2}.$
Since by Lemma 19, $\inf_{\widehat{V}}\sup_{V\in\mathcal{V}}\mathbb{E}\|\widehat{V}-V\|^{2}\geq c_{2}\kappa^{2}d,$ it follows that
$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{B}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}=\inf_{\widehat{V}}\sup_{V\in\mathcal{V}}\mathbb{E}\|\widehat{V}-V\|^{2}\geq c_{2}\kappa^{2}d=c_{2}\kappa^{2}(r_{1}+r_{1}^{\prime})r_{2}r_{3}.$
Step 1.2. To show (27), there are two cases. (a) Suppose $\xi^{2}\geq r_{2}r_{3}\kappa^{2}$, then $r_{1}^{\prime}=\lfloor\xi^{2}/(r_{2}r_{3}\kappa^{2})\rfloor\geq 1$ and therefore
$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}}\mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\geq c_{2}\kappa^{2}(r_{1}+r_{1}^{\prime})r_{2}r_{3}\geq c_{2}r_{1}r_{2}r_{3}+c_{3}\xi^{2}.$
(b) Suppose $\xi^{2}<r_{2}r_{3}\kappa^{2}$. Then $r^{\prime}_{1}=0$
$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\$$\geq\ c_{2}\kappa^{2}r_{1}r_{2}r_{3}\geq\frac{c_{2}}{2}(\kappa^{2}r_{1}r_{2}r_{3}+\xi^{2}),$
where the first inequality follows from Step 1.1, the second inequality follows from $\xi^{2}<r_{2}r_{3}\kappa^{2}$.
Step 2. Note that
$\inf_{\widehat{X}}\ \sup_{X^{*}\in\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\geq\inf_{\widehat{X}}\ \sup_{X^{*}\in\mathcal{T}_{(r_{1},r_{2},r_{3})}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\ \geq\ c_{3}\kappa^{2}\sum_{k=1}^{3}p_{k}r_{k},$ (30)
where the second inequality follows from Zhang and Xia (2018). The desired result directly follows from (27) and (30). ∎
Appendix C SVD for Unbalanced Matrices
@@
where the last inequality uses $\frac{\sigma_{R}^{2}(X)+n{\kappa}^{2}}{n\,\kappa^{2}}\geq 1$. Thus, combining (C), (44), (45) and (46) with $x=\kappa\sqrt{\frac{2C}{c}m\,}$, we have
@@
It follows from (C), (42) and (47) that by the union bound,
@@
Suppose the conditions of Theorem 3 hold, in particular the condition in Equation (6) for each mode-$k$ unfolding $\mathcal{M}_{k}(X^{*})$ with target rank $r_{k}$. Let $U_{k}^{(0)}$ be the matrix of top $r_{k}$ left singular vectors of the mode-$k$ unfolding $Y^{(k)}=\mathcal{M}_{k}(Y)$. Then for each $k\in\{1,2,3\}$, with probability at least $1-C_{1}\exp(-C_{2}p_{k})$,
@@
Apply Lemma 1 to $Y^{(k)}=\mathcal{M}_{k}(Y)\in\mathbb{R}^{p_{k}\times p_{\neq k}}$ with signal $X^{(k)}=\mathcal{M}_{k}(X^{*})$ and noise $Z^{(k)}=\mathcal{M}_{k}(Z)$, where $p_{\neq k}=\prod_{j\neq k}p_{j}$. Use rank $r=r_{k}$ and identify $n=p_{k}$ (rows) and $m=p_{\neq k}$ (columns). The lemma (applied to left singular vectors, or equivalently to the transpose) gives
$\bigl\|\sin\Theta(U_{k}^{(0)},U_{k}^{*})\bigr\|^{2}\;\leq\;C\left\{\frac{p_{k}\,\kappa^{2}}{\Delta_{k}^{2}}+\frac{\kappa^{4}\,p_{k}\,p_{\neq k}}{\Delta_{k}^{4}}\right\},$
where $\Delta_{k}=\sigma_{r_{k}}\!\bigl(\mathcal{M}_{k}(X^{*})\bigr)-\sigma_{r_{k}+1}\!\bigl(\mathcal{M}_{k}(X^{*})\bigr)$ is the mode-$k$ spectral gap.
@@
$\Delta_{k}^{2}\;\geq\;C_{\mathrm{gap}}\kappa^{2}\!\left(\sqrt{p_{k}\,p_{\neq k}\,r_{\max}}+r_{\max}\sum_{j=1}^{3}p_{j}\right).$
@@
Appendix D Auxiliary results
D.1 Deviations bounds
Lemma 2 (Theorem 4.4.3 in Vershynin (2018)).
Then there exist some universal constant $C>0$, such that
$\|Z\|\leq C\kappa\left(\sqrt{m}+\sqrt{n}\right)$
with probability at least $1-\exp(-(m+n)))$.
Lemma 3.
Then there exist some universal constants $C,c>0$, such that
with probability at least $1-\exp(-c\sum_{k=1}^{3}p_{k}r_{k})$.
Proof.
This directly follows from Lemma E.5 in Han et al. (2022). ∎
Lemma 4.
Proof.
Lemma 5.
Then there exists absolute positive constants $C_{1},C_{2}$ and $c$ such that
$\leq C_{2}\exp\left(-c(p_{1}+p_{2}+p_{3})\right),$
Proof.
It follows from Lemma 6 that
$\{A\in\mathbb{R}^{p_{2}\times r_{2}}:\|A\|\leq 1\}$
$\{B\in\mathbb{R}^{p_{3}\times r_{3}}:\|B\|\leq 1\}$
Then
Note that $\|V_{2}-\widetilde{V}_{2}\|\leq 1/4$. So
Similarly
In addition,
So for any $V_{2}$ and $V_{3}$,
or simply
Step 3. By Proposition 8 in Pajor (1998), the cardinality of $\mathcal{N}_{p_{2},r_{2}}(\epsilon)$ is bounded $(\frac{C}{\epsilon})^{p_{2}r_{2}}$, and $\mathcal{N}_{p_{3},r_{3}}(\epsilon)$ is bounded $(\frac{C}{\epsilon})^{p_{3}r_{3}}$. Therefore
Lemma 6.
Proof.
$\|a-\widetilde{a}\|_{2}\leq\epsilon.$
Denote the random quantity
Then
[…]
Note that $\|a-\widetilde{a}\|\leq\frac{1}{4}$. So
Similarly
$|\widetilde{a}^{\top}AZB(\widetilde{b}-b)|\leq\frac{\psi}{4}.$
In addition,
So for any $a$ and $b$,
or simply
Step 4. By Vershynin (2018), the cardinality of $\mathcal{N}_{p}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{p}$, and the cardinality of $\mathcal{N}_{q}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{q}$. Therefore
Lemma 7.
where $C$ and $C_{1}$ are positive constants.
Proof.
For any non-random $u,v\in\mathbb{R}^{m}$ , it follows that
$\|a-\widetilde{a}\|_{2}\leq\epsilon.$
Denote the random quantity
Then
[…]
Note that $\|a-\widetilde{a}\|_{2}\leq\frac{1}{4}$. So
Similarly
In addition,
So for any $a$ and $b$,
or simply
Step 2. By Vershynin (2018), the cardinality of $\mathcal{N}_{p}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{p}$, and the cardinality of $\mathcal{N}_{q}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{q}$, for a positive constant $C$. Therefore
Lemma 8.
$\|u^{\top}Zv\|_{\psi_{2}}\leq\kappa\|u\|_{2}\|v\|_{2}.$
Then with probability at least $1-\exp(-c(m+n))$, it holds that
where $c,C$ and $C_{1}$ are positive constants.
Proof.
By assumption,
The rest of the proof is similar and simpler than Lemma 6 and is omitted.
D.2 Matrix perturbation bounds
@@
where $B_{(R)}$ denote the rank-$R$ truncated SVD of $B$, this is, if $B=U\Sigma V^{\top}$ then $B_{(R)}=U_{(R)}\Sigma_{(R)}V_{(R)}^{\top}$.
@@
where the first inequality follows from the triangle inequality, and second inequality follows from Weyl’s inequality (Weyl, 1912), i.e. $|\sigma_{j}(A)-\sigma_{j}(B)|\leq\|A-B\|$ for all $1\leq j\leq n$, as well as the fact that
@@
where the last inequality follows from Theorem 5. It then follows from (D.2),
@@
where the first two inequalities follow from the same arguments as in (D.2). Consequently,
@@
Lemma 15 (Proposition 1 of Cai and Zhang (2018)).
@@
Theorem 5 (Mirsky’s singular value inequality in Mirsky (1960)).
@@
Lemma 18 (Ky Fan-type Inequality for Sums of Matrices).
Proof.
For a symmetric matrix $M\in\mathbb{R}^{n\times n}$, by Ky Fan’s maximum principle (see e.g. II.1.13 in Bhatia (2013)), for any $1\leq k\leq n$,
Therefore
and so
Then
D.3 Minimax lower bound for vectors
Lemma 19.
Let $\kappa>0$ and $d\in\mathbb{Z}^{+}$, and denote
$\mathcal{V}=\{V\in\mathbb{R}^{d},\|V\|_{\infty}\leq\kappa\}.$
Consider the model $Y\sim\mathcal{N}(V,\kappa^{2}I_{d})$ where $V\in\mathcal{V}$. Then there exists a universal constant $c$ such that
$\inf_{\widehat{V}}\sup_{V\in\mathcal{V}}\mathbb{E}\|\widehat{V}-V\|^{2}\geq c\kappa^{2}d,$
where infimum is taken over all the estimators $^{\widehat{V}}$ based on $Y$.
Proof.
This is a standard minimax lower bound for bounded normal means; see for example Berry (1990) and Donoho et al. (1990). ∎